A person is pushing a sled along a horizontal surface. They exert a force of 200 N on the 75 kg sled. In addition a frictional force of 20 N acts on the sled.

What is the rate of acceleration of the sled? (in m s−2 to 2.s.f)


1
Expert's answer
2021-04-19T17:10:42-0400

Solution.

"F=200N;"

"m=75kg;"

"F_f=20N;"

"a-?;"

"\overrightarrow{ma}=\overrightarrow{F}+\overrightarrow{F_f};"

"ma=F-F_f\implies a=\dfrac{F-F_f}{m};"

"a=\dfrac{200N-20N}{75kg}=2.40ms^{-2};"

Answer: "a=2.40ms^{-2}."